OpenStudy (anonymous):

Consider the two vectors A and B and their resultant A+B. The magnitudes of the vectors A and B are, respectively 18.8 and 7.6 and they act 100 degrees to each other. Find the magnitude of the resultant vector A+B Find the angle between the direction of the resultant vector A+B and the direction of the vector A I know we use the law of cosine but I keep getting it wrong :/

OpenStudy (anonymous):

 What did u get ?

OpenStudy (***[isuru]***):

Hi, The resultant (R)^2 = A^2 + B^2 + 2ABcos(x) let (x) be 100 then R^2 = (18.8)^2 + (7.6)^2 +(2*18.8*7.6* cos 100) = 460.80 R= 21.46 And @oksuz_ why did u use 80 degree instead of 100?

OpenStudy (anonymous):

if we want to find A-B, the resultant vector become like this  100 degree angle sees this resultant vector, so according to law of cosine we have to arrange the angle what is seen by resultant vector.